OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (luigi0210):

Solve the second equation for y: y=-4x-1 Then plug it in to the other equation: 6x+2(-4x-1)=0 6x-8x-2=0 -2x=2 x=-1 Plug in x and solve for y. 4x+y=-1 4(-1)+y=-1 -4+y=-1 y=3